OpenStudy (anonymous):

graph the inequality. 4<6t+1<=43

OpenStudy (saifoo.khan):

6 < 6t + 1 6t + 1 <= 43 Solve both.

OpenStudy (anonymous):

when you divide do you switch the less then greater than signes?

OpenStudy (saifoo.khan):

no. just make t the subject for both. then combine.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

well the first one is t<1.2

OpenStudy (saifoo.khan):

6 < 6t + 1 6t + 1 <= 43 5/6 < t t <= 42 5/6 < t < 42

OpenStudy (anonymous):

so then cimbine 1.2 +45?

OpenStudy (anonymous):

errr 42

OpenStudy (saifoo.khan):

nope, that's it

OpenStudy (anonymous):

how would u graph that ?

OpenStudy (anonymous):

dont u have to divide 42 by 6?

OpenStudy (saifoo.khan):

omg! i did two blunders

OpenStudy (saifoo.khan):

sorry.

OpenStudy (anonymous):

its okay lol im still confused tho. So you divide by 6 in both equations and then add the like terms together right ?

OpenStudy (saifoo.khan):

4<6t+1<=43 4 < 6t + 1 6t + 1 <= 43 3 < 6t 6t <= 42 0.5 < t t <= 7 0.5 < t <= 7

OpenStudy (anonymous):

so then how do u graph that ? ive never seen inequalities like this

OpenStudy (saifoo.khan):